OpenStudy (anonymous):

A line crosses the y-axis at (0,4) and has a slope of -2. Find an equation for the this line. A) y = 2x + 4 B) y = -2x + 4 C) y = -2x - 4 D) y = -4x + 2

OpenStudy (texaschic101):

y = mx + b slope(m) = -2 (0,4)....x = 0 and y = 4 now we sub 4 = -2(0) + b 4 = b so your equation is : y = -2x + 4

OpenStudy (anonymous):

@jwhite12 obviously, but i don't know what numbers go where or i wouldn't be asking. @texaschic101 thank you

OpenStudy (texaschic101):

sure thing :)